# QueSTMaps: Queryable Semantic Topological Maps for 3D Scene Understanding

### v1

@@ Abstract
Understanding the structural organisation of 3D indoor scenes in terms of rooms is often accomplished via floorplan extraction. Robotic tasks such as planning and navigation require a semantic understanding of the scene as well. This is typically achieved via object-level semantic segmentation. However, such methods struggle to segment out topological regions like “kitchen” in the scene. In this work, we introduce a two-step pipeline. First, we extract a topological map, i.e., floorplan of the indoor scene using a novel multi-channel occupancy representation. Then, we generate CLIP-aligned features and semantic labels for every room instance based on the objects it contains using a self-attention transformer. Our language-topology alignment supports natural language querying, e.g., a “place to cook” locates the “kitchen”. We outperform the current state-of-the-art on room segmentation by $\sim$20and room classification by $\sim$12%. Our detailed qualitative analysis and ablation studies provide insights into the problem of joint structural and semantic 3D scene understanding.
Figure 1: Topology Extraction and Room Labeling for Indoor Scene 3D Point clouds. Given a 3D scene point cloud we (a) predict room and transition regions, (b) generate room-label aligned CLIP embeddings to assign room labels, and (c) build a topological map that supports room-level natural language queries, e.g., the query place to cook locates the kitchen.
@@ I Introduction
Understanding both the structure and the semantics of 3D indoor environments is essential for mapping and navigation in real world [1, 2, 3, 4]. The structure, i.e., the layout or floorplan, of an indoor environment defines its organisation into different rooms or spaces. Each room has a semantic label associated with it, e.g., kitchen, which can be inferred from the semantic labels of the objects inside it. Thus, we can generate a topological understanding of such 3D scenes using connections both within and across the rooms, that is, objects-to-room and room-to-room.
Very few existing works address the twin problem of room segmentation and classification, with most focusing on segmentation alone. Floorplan extraction methods attempting to understand only the structure of the scene [5], ignore the rich semantics essential for high-level robotic tasks [6, 7]. On the other hand, dense 3D semantic segmentation methods [8, 9, 10, 11] have focused solely on object-level semantic segmentation, without considering the room-level structure of indoor scenes [12].
In this paper, we propose a first-of-its-kind method that generates a natural language-queryable topological and semantic representation of a 3D indoor scene. We propose a novel multi-channel occupancy representation of 3D point clouds for room segmentation. These channels help in disambiguating the transition regions (e.g., doors) and rooms in the scene. We generate room-level features for each room segment using the CLIP [13] embeddings of objects confined in them. We align the room features with natural language descriptions of the room using a self-attention transformer. As a result, our method supports free-form natural language queries to detect relevant rooms. For example, the query a place to cook locates the kitchen. We benchmark our method on synthetic datasets like Structured3D [14] and challenging real-world datasets like Matterport3D [15]. We achieve $\sim$20and $\sim$12performance gain over the previous state-of-the-art methods on floorplan extraction [16, 5] and room labeling [17] respectively. Our research contributions are listed below:
@@ II-A Floorplan Reconstruction Methods
Floorplan reconstruction organises raw sensor data like point clouds, density maps, or RGB-D images into a structured representation of rooms. Classical methods leverage techniques like plane-fitting [18, 19] or model the problem as an energy minimisation problem [20, 21]. The heuristic-based detection of such methods struggles with noisy and incomplete real world 3D data. Deep learning based methods tackle this concern [22, 23, 24], where some methods [25, 26] such as HEAT [5] use transformers to generate point, edge, or room level proposals to generate a floorplan. These methods do not predict semantics of the indoor environment (i.e., no room labels like bedroom, kitchen, etc.) and are thus limited in their downstream utility [3, 27]. Furthermore, unlike aforementioned works, in this paper, we also seek to identify transition regions, which not only help in room segmentation but can also aid in downstream navigation [28].
@@ II-B Semantic Segmentation of Floorplans
Semantic segmentation of floorplans adds the challenge of assigning a semantic label to every detected room [29]. This often includes transition regions like doorways and windows. FloorNet [24] utilises a CNN-DNN architecture and also identifies salient objects in the scene. However, its object set is very limited. RoomFormer [16], the current state-of-the-art, models the problem as a polygon fitting task by regressing to vertices and edges. Although such a formulation suits synthetic scenes [14] with straight edges and regular angles, it struggles on real-world scenes [15] which invariably contain noise, ill defined edges, and curved walls. These methods also struggle to segment rooms with cluttered objects. Since these methods do not perform dense semantic segmentation, they do not have open vocabularies and cannot be queried in natural language.
@@
3D semantic segmentation methods assign a label to 3D pixels or voxels in the scene [30]. Methods like OpenScene [8, 9, 10, 11] use per-pixel image features extracted from posed RGB-D images of a scene and obtain a point-wise task-agnostic scene representation but struggle in differentiating instances of objects. 3D instance segmentation methods [12, 31, 32, 33] distinguish multiple objects belonging to the same semantic category, predicting individual masks for each instance, but have no understanding of how the objects are structurally associated. Thus, they cannot label entire regions like kitchens, bedrooms etc.
II-D 3D Semantic Scene Graphs
To unite structure with semantics [34, 35, 36, 37] seek to build 3D scene graphs where nodes represent objects and edges encode inter-object relationships. To perform downstream tasks we must be able to query these graphs conveniently. ConceptGraphs [38] and others [39, 40, 41, 42] use LLMs to accomplish this due to their open-vocabulary and reasoning abilities. [43] incrementally grounds spaces conditioned on language. However, all these methods struggle to segment rooms from unstructured data, and cannot retrieve all objects from a queried room.
II-E 3D Scene Understanding
For a robot to understand a scene it must combine place and object labels in a hierarchical manner to create a hybrid map [7][44] jointly classifies objects and room labels, and can identify novel objects, but its model parameters such as room adjacency and room-object compatibility are manually assigned, limiting its generality. [45] uses an unsupervised geometry-based approach to segment 3D point clouds into objects and meaningful scene structures but its segmentations lack semantics. [46] can recognize new semantic classes online, but does not leverage rich 3D object geometries and is limited to semantic segmentation in 2D space. [47] presents TopoNets for semantic mapping based on the topological structure of the environment, but lacks an open vocabulary. Although the aforementioned techniques can segment an environment topologically into meaningful spaces, they are not natural language-queryable.
@@ III-A1 Multi-Channel Occupancy Representation
$O_{ij}=\begin{cases}1&\text{if }N_{ij}\geq 0\\ 0&\text{otherwise}\end{cases}$ (1)
$N_{ij}=\sum_{p\in\mathcal{P}}\mathbb{I}(\beta_{1}h\leq p_{z}\leq\beta_{2}h)$ (2)
where $O_{ij}$ and $N_{ij}$ are respectively the occupancy state and count of the corresponding grid cell, and $\beta_{1}$ and $\beta_{2}$ define the extent of the slice. For the ceiling slice, we choose $\beta_{1}=0.7$ and $\beta_{2}=0.9$, and for the floor slice, we choose $\beta_{1}=0.1$ and $\beta_{2}=0.3$. $p$ represents a point in the point cloud, $p_{z}$ is the z-coordinate of the point, $h$ is the height of the point cloud calculated as the difference between the ceiling and the floor heights, and $\mathbb{I}(\cdot)$ is the indicator function.
As illustrated in Figure Figure 2, these additional input channels clearly delineate doorways as a difference between the floor and ceiling occupancy maps, simplifying the problem of doorway/transition region detection, and consequently room segmentation.
@@ III-A2 Room-Mask Prediction and Transition Region Detection
We formulate room instance mask prediction and transition region detection as a 3-class (2 classes + background) instance segmentation task, similar to Mask R-CNN. The input for this segmentation task is our three-channel 2D array consisting of the density map and the top and bottom occupancy maps, as detailed in Sec. subsubsection III-A1. The total loss in Mask R-CNN is given by:
@@ III-B Room Labeling
We introduce a novel approach to assign room labels to the room masks generated in Sec. subsection III-A. To achieve this, we combine these room segments with their confined objects, where we obtain object instances using a 3D semantic instance mapping pipeline [38]. Using the semantic information for the objects in a room, we use our model to obtain room embeddings aligned with the CLIP features of room labels.
@@
We use the semantic object map creation methodology described in ConceptGraphs [38] to obtain a dense object instance map, as detailed in this sub-subsection for the sake of completion. This map is derived from a series of posed RGB-D images, collectively represented as $I=\{I_{1},I_{2},\ldots,I_{t}\};I_{t}=\langle I_{t}^{rgb},I_{t}^{depth},\theta_{t}\rangle$. It lifts the outputs of 2D foundation models into 3D by aggregating observations associated across multiple views. This results in distinct point clouds and features for each object. Segment Anything Model (SAM) [48] is used to obtain class-agnostic instance masks within each image, denoted as $\{m_{t,i}\}_{i=1}^{M}=\text{Seg}(I_{t}^{rgb})$. The corresponding mask features $f_{t,i}=\text{Embed}(I_{t}^{rgb},m_{t,i})$ are computed using CLIP [13] by cropping the masked image. Each object within the scene, $o_{j}$, is thus defined by a tuple $o_{j}=\langle p_{o_{j}},f_{o_{j}}\rangle$, where $p_{o_{j}}$ represents the object’s point cloud and $f_{o_{j}}$ represents the associated CLIP embedding for the object. This point cloud $p_{o_{j}}$ fuses the constituent point clouds $p_{i}$ through multi-view projection of its 2D masks $m_{t,i}$ into 3D using depth $I_{t}^{depth}$ and pose $\theta_{t}$. In doing so, CLIP features $f_{t,i}$ are averaged across views to obtain a single object CLIP feature vector $f_{o_{j}}\in\mathbb{R}^{D}$. This feature encapsulates the semantic attributes of the object.
@@ III-B2 Object-to-Room Association
After obtaining the room segmentation masks $R=\{r_{1},r_{2},\ldots,r_{n}\}$ as detailed in Sec. subsection III-A, we allocate object instances to specific rooms. We first find out which storey of the indoor scene the object is on. We then apply a transformation to the segmentation masks to convert from 2D image space back to 3D world co-ordinates. This transformation is facilitated by the offset and tile-size metadata we save during the point cloud pre-processing step detailed in subsection subsubsection III-A1
We employ a KD Tree construction for each room mask, incorporating all $(x,y)$ coordinates as defined in the world frame. This indexing enables efficient querying of the room mask and rapid association of objects to rooms. For each object identified within the scene, we compute the centroid of its point cloud representation. This centroid serves as a query point against the constructed KD Trees, with the closest match dictating the room assignment for the respective object. Utilizing this technique, we are able to accurately relate discrete object instances to their corresponding room masks.
@@ III-B3 Assigning Room Labels
We aim to maximize the similarity between the CLS token embedding and the ground truth room label’s text embedding using a contrastive loss function, similar to the Normalized Temperature-scaled Cross Entropy Loss (NT-Xent). This refines the model’s ability to output a room label aligned embedding $e_{CLS}$ , which is in the CLIP embedding space.
@@
Additionally, the output embedding is designated as the room embedding for the associated 3D region. This facilitates a targeted search for specific room types using natural language queries, since the room embedding is aligned with CLIP.
@@ III-C Training Details
Our self-attention transformer model for room labeling incorporates an embedding size of 1024, with 8 attention heads, and 8 layers. During training, we set the dropout at a rate of 0.2 to prevent overfitting and the temperature $\tau$ for the loss function to 0.5. Optimization is performed using AdamW, with a learning rate of $1\times 10^{-5}$ and training for 400 epochs.
@@ IV Experiments and Results
Sec. subsection IV-A discusses datasets, metrics, ablations, results, and comparisons of the first part of our hierarchical segmentation pipeline, i.e. room segmentation and transition region detection. Sec. subsection IV-B discusses labeling rooms and finding room labels for each room mask. In Sec. subsection IV-C we evaluate the complete pipeline and benchmark it against RoomFormer [16] on Matterport3D [15]. We also show qualitative comparisons with other open-vocabulary 3D scene understanding methods to highlight our method’s ability to understand the topology of 3D scenes.
@@ IV-A Room and Transition Region Segmentation
In this part of the topological segmentation pipeline, we train a model based on the Mask R-CNN architecture [22] to segment a scene into instance-separated rooms as well as instance-separated doors, i.e. transitions between the rooms.
Datasets. We evaluate our method on two datasets-Structured3D and Matterport3D [14, 15]. Structured3D comprises panoramic Manhattan and non-Manhattan floorplan layouts with semantically rich room annotations. Its scenes are synthetic with minimal clutter. Notably, the rooms within this dataset are geometrically defined as polygons, a feature that has facilitated the efficacy of previous works  [16, 5] based on polygon fitting techniques. We adhere to the official split of 3000 training, 250 validation, and 250 test samples. Matterport3D provides RGBD scans and navigable meshes of real-world indoor scenes with object and room type annotations. It is used extensively for robotic tasks like navigation and 3D scene understanding. It offers very detailed scenes with cluttered rooms. The major challenges associated with it are the smaller number of scenes (90), class imbalance in room types, and noisy meshes. Moreover, it also contains open-plan scenes and open doors. Although Matterport3D is a popular dataset for 3D indoor scene understanding, no other work discusses topological segmentation of scenes for this dataset. We split multistorey scenes using K-means clustering over the z-coordinates of room centers to obtain a distinct point cloud for each storey. Each storey’s point cloud is clipped at an estimated ceiling and floor heights found by running a 1-D peak-finding algorithm on the histogram of z-coordinates of the point cloud. Since Matterport3D does not provide annotations for transition regions, we manually annotated all doorways in the dataset. We share this extension to the dataset with the community.
@@
Figure 3: Qualitative comparisons on room segmentation task. Room segmentations produced by Roomformer and our method on Structured3D and Matterport3D scenes. Roomformer fails in cases where rooms have curved walls and clutter (fourth row). In cases where the density is low, Roomformer fails to predict vertices which adversely affects the segmentation (second row). Colors are only representative of the instance separation.
Baselines and Results. We primarily compare our model against the state-of-the-art RoomFormer [16] and HEAT [5]. Both are transformer-based methods that fit polygons and output an ordered set of vertices. [5] first detects room corners, and then predicts edges, but no doorways. Results are summarized in Table Table I, where we outperform the average precision of the state-of-the-art RoomFormer by 1.9% on Structured3D and by 20% on Matterport3D. In addition to rooms, RoomFormer also detects transition regions i.e. doorways. Our method registers a 10% higher average precision than the state-of-the-art RoomFormer on doorway detection. To demonstrate the effectiveness of our multi-channel representation (top and bottom occupancy maps), we construct a baseline that learns only on a single-channel density map. As in Table Table I, our ablations exhibit an improvement in both doorway detection and room segmentation tasks.
Figure 4: Topological Map Generation. From the outputs of the room segmentation and room labeling pipelines, we build topological maps where nodes represent both rooms as well as transition regions and edges signify connectivity. This representation enables downstream tasks such as topology informed language commands in a robot navigation setting.

### v2

@@ Abstract
Robotic tasks such as planning and navigation require a hierarchical semantic understanding of a scene, which could include multiple floors and rooms. Current methods primarily focus on object segmentation for 3D scene understanding. However, such methods struggle to segment out topological regions like “kitchen” in the scene. In this work, we introduce a two-step pipeline to solve this problem. First, we extract a topological map, i.e., floorplan of the indoor scene using a novel multi-channel occupancy representation. Then, we generate CLIP-aligned features and semantic labels for every room instance based on the objects it contains using a self-attention transformer. Our language-topology alignment supports natural language querying, e.g., a “place to cook” locates the “kitchen”. We outperform the current state-of-the-art on room segmentation by $\sim$20and room classification by $\sim$12%. Our detailed qualitative analysis and ablation studies provide insights into the problem of joint structural and semantic 3D scene understanding. Project Page: quest-maps.github.io
[Uncaptioned image] Figure 1: Topology Extraction and Room Labeling for Indoor Scene 3D Point clouds. Given a 3D scene point cloud we (a) predict room and transition regions, (b) generate room-label aligned CLIP embeddings to assign room labels, and (c) build a topological map that supports room-level natural language queries, e.g., the query place to cook locates the kitchen.
@@ I Introduction
Understanding both the structure and the semantics of 3D indoor environments is essential for mapping and navigation in real world [1, 2, 3]. The structure, i.e., the layout or floorplan, of an indoor environment defines its organisation into different rooms or spaces. Each room has a semantic label associated with it, e.g., kitchen, which can be inferred from the semantic labels of the objects inside it. Thus, we can generate a topological understanding of such 3D scenes using connections both within and across the rooms, that is, objects-to-room and room-to-room. Such a topological understanding coupled with a semantic understanding helps plan paths across several rooms, when given a free-form natural language query.
Very few existing works address the twin problem of room segmentation and classification, with most focusing on segmentation alone. Floorplan extraction methods attempting to understand only the structure of the scene [4], ignore the rich semantics essential for high-level robotic tasks [5, 6]. On the other hand, dense 3D semantic segmentation methods [7, 8] have focused solely on object-level semantic segmentation, without considering the room-level structure of indoor scenes [9].
We propose a first-of-its-kind method that generates a natural language-queryable room-level topological and semantic representation of a 3D indoor scene, by segmenting then understanding. We propose a novel multi-channel occupancy representation of 3D point clouds for room segmentation. These channels help in disambiguating the transition regions (e.g., doors) and rooms. We generate room-level features for each room segment using the CLIP [10] embeddings of objects confined in them. We align the room features with natural language descriptions of the room using a self-attention transformer. As a result, our method supports free-form natural language queries to detect relevant rooms. For example, the query a place to cook locates the kitchen. We benchmark our method on synthetic datasets like Structured3D [11] and challenging real-world datasets like Matterport3D [12]. We achieve $\sim$20and $\sim$12performance gain over the previous SoTA methods on floorplan extraction [13, 4] and room labeling [14] respectively. Our contributions are:
@@ II-A Floorplan Reconstruction Methods
Floorplan reconstruction organises raw sensor data like point clouds, density maps, or RGB-D images into a structured representation of rooms. Classical methods leverage techniques like plane-fitting [15] or model the problem as an energy minimisation problem [16]. The heuristic-based detection of such methods struggles with noisy and incomplete real world 3D data. Deep learning based methods tackle this concern [17, 18, 19], where some methods [20, 21] such as HEAT [4] use transformers to generate point, edge, or room level proposals to generate a floorplan. These methods do not predict semantics of the indoor environment (i.e., no room labels like bedroom, kitchen, etc.) and are thus limited in their downstream utility [2].
@@ II-B Semantic Segmentation of Floorplans
This includes assigning a semantic label to every detected room, and transition regions like doorways and windows. FloorNet [19] utilises a CNN-DNN architecture and also identifies salient objects in the scene. However, its object set is very limited. RoomFormer [13], the current state-of-the-art, models the problem as a polygon fitting task by regressing to vertices and edges. Although such a formulation suits synthetic scenes [11] with straight edges and regular angles, it struggles on real-world scenes [12] which invariably contain noise, ill defined edges, and curved walls. These methods also struggle to segment rooms with cluttered objects. Since these methods do not perform dense semantic segmentation, they do not have open vocabularies and cannot be queried in natural language.
@@
3D semantic segmentation methods assign a label to 3D pixels or voxels in the scene [22]. Methods like OpenScene [7, 8] use per-pixel image features extracted from posed RGB-D images of a scene and obtain a point-wise task-agnostic scene representation but struggle in differentiating instances of objects. 3D instance segmentation methods [9, 23, 24] distinguish multiple objects belonging to the same semantic category, predicting individual masks for each instance, but do not build structural associations among objects. Thus, they cannot label entire regions like kitchens, bedrooms etc.
II-D 3D Scene Understanding and Scene Graphs
For a robot to understand a scene it must combine place and object labels in a hierarchical manner to create a hybrid map [6][25] uses an unsupervised geometry-based approach to segment 3D point clouds into objects and meaningful scene structures but its segmentations lack semantics. [26] can recognize new semantic classes online, but does not leverage rich 3D object geometries and is limited to semantic segmentation in 2D space. To unite structure with semantics [27, 28, 29, 30] seek to build 3D scene graphs where nodes represent objects and edges encode inter-object relationships. To perform downstream tasks we must be able to query these graphs conveniently. ConceptGraphs and others [31, 32, 33, 34] use LLMs to accomplish this due to their open-vocabulary and reasoning abilities. However, all these methods struggle to segment rooms from unstructured data, and cannot retrieve all objects from a queried room.
@@ III-A1 Multi-Channel Occupancy Representation
$O_{ij}=\mathbb{I}\left(\sum_{p\in\mathcal{P}}\mathbb{I}(\beta_{1}h\leq p_{z}\leq\beta_{2}h)\geq 0\right)$ (1)
where $O_{ij}$ is the occupancy state of the grid cell, and $\beta_{1}$ and $\beta_{2}$ define the extent of the slice. For the ceiling slice, we choose $\beta_{1}=0.7$ and $\beta_{2}=0.9$, and for the floor slice, we choose $\beta_{1}=0.1$ and $\beta_{2}=0.3$. $p$ represents a point in the point cloud, $p_{z}$ is the z-coordinate of the point, $h$ is the height of the point cloud calculated as the difference between the ceiling and the floor heights, and $\mathbb{I}(\cdot)$ is the indicator function.
As illustrated in Fig. Figure 2, these additional input channels clearly delineate open doorways as a difference between the floor and ceiling occupancy maps, simplifying the problem of doorway/transition region detection, and consequently room segmentation.
@@ III-A2 Room-Mask Prediction and Transition Region Detection
We formulate room instance mask prediction and transition region detection as a 3-class (2 classes + background) instance segmentation task. The input for this segmentation task is our three-channel 2D array consisting of the density map and the top and bottom occupancy maps, as detailed in Sec. subsubsection III-A1. The total loss in Mask R-CNN is given by:
@@ III-B Room Labeling
We introduce a novel approach to assign room labels to the room masks generated in Sec. subsection III-A. To achieve this, we combine these room segments with their contained objects, where we obtain object instances using a 3D semantic instance mapping pipeline [31]. Using the semantic information for the objects in a room, we use our model to obtain room embeddings aligned with the CLIP features of room labels.
@@
We use the semantic object map creation methodology described in ConceptGraphs [31] to obtain a dense object instance map, as detailed in this sub-subsection for the sake of completion. This map is derived from a series of posed RGB-D images, collectively represented as $I=\{I_{1},I_{2},\ldots,I_{n}\};I_{t}=\langle I_{t}^{rgb},I_{t}^{depth},\theta_{t}\rangle$. For some $1\leq t\leq n$. It lifts the outputs of 2D foundation models into 3D by aggregating observations associated across multiple views. This results in distinct point clouds and features for each object. Segment Anything Model (SAM) [35] is used to obtain class-agnostic instance masks within each image, denoted as $\{m_{t,i}\}_{i=1}^{M}=\text{Seg}(I_{t}^{rgb})$. The corresponding mask features $f_{t,i}=\text{Embed}(I_{t}^{rgb},m_{t,i})$ are computed using CLIP [10] by cropping the masked image. Each object within the scene, $o_{j}$, is thus defined by a tuple $o_{j}=\langle p_{o_{j}},f_{o_{j}}\rangle$, where $p_{o_{j}}$ represents the object’s point cloud and $f_{o_{j}}$ represents the associated CLIP embedding for the object. This point cloud $p_{o_{j}}$ fuses the constituent point clouds $p_{i}$ through multi-view projection of its 2D masks $m_{t,i}$ into 3D using depth $I_{t}^{depth}$ and pose $\theta_{t}$. In doing so, CLIP features $f_{t,i}$ are averaged across views to obtain a single object CLIP feature vector $f_{o_{j}}\in\mathbb{R}^{D},D=1024$. This feature encapsulates the semantic attributes of the object.
@@ III-B2 Object-to-Room Association
After obtaining the room segmentation masks $R=\{r_{1},r_{2},\ldots,r_{n}\}$ as detailed in Sec. subsection III-A, we allocate object instances to specific rooms. We apply a transformation to the segmentation masks to convert from 2D image space back to 3D world co-ordinates. This transformation is facilitated by the offset and tile-size metadata we save during the point cloud pre-processing step detailed in subsection subsubsection III-A1
We employ a KD Tree construction for each room mask, to accurately relate discrete object instances to their corresponding room masks. incorporating all $(x,y)$ coordinates as defined in the world frame. For each object identified within the scene, we compute the centroid of its point cloud. This centroid serves as a query point against the constructed KD Trees, with the closest match dictating the room assignment for the respective object.
@@ III-B3 Assigning Room Labels
We aim to maximize the similarity between the CLS token embedding and the ground truth room label’s text embedding using a contrastive loss function, similar to the Normalized Temperature-scaled Cross Entropy Loss. This refines the model’s ability to output a room label aligned embedding $e_{CLS}$ , which is in the CLIP embedding space.
@@
Additionally, the output is designated as the room embedding for the associated 3D region. This facilitates a targeted search for specific room types using natural language queries, since the room embedding is aligned with CLIP.
@@ III-C Training Details
Our self-attention transformer model for room labeling incorporates an embedding size of 1024, with 8 attention heads, and 8 layers. During training, we set the dropout at a rate of 0.2 to prevent overfitting and the temperature $\tau$ for the loss function to 0.5. Optimization is performed using AdamW, with a learning rate of $1\times 10^{-5}$ for 400 epochs.
@@ IV Experiments and Results
Sec. subsection IV-A discusses datasets, metrics, ablations, results, and comparisons of the room segmentation and transition region detection step. Sec. subsection IV-B discusses labeling rooms and finding room labels for each room mask. In Sec. subsection IV-C we evaluate the complete pipeline and benchmark it. We also show qualitative comparisons to highlight our method’s ability to understand the topology of 3D scenes.
@@ IV-A Room and Transition Region Segmentation
This part of the topological segmentation pipeline, utilises a MaskRCNN  [17] to segment a scene into instance-separated rooms as well as instance-separated doors, i.e. transitions between the rooms.
Datasets. We evaluate our method on two datasets-Structured3D and Matterport3D [11, 12]. Structured3D comprises panoramic Manhattan and non-Manhattan floorplan layouts with semantically rich room annotations with synthetic scenes with minimal clutter. Notably, the rooms within this dataset are geometrically defined as polygons, a feature that has facilitated the efficacy of previous works  [13, 4] based on polygon fitting techniques. We adhere to the official split of 3000 training, 250 validation, and 250 test samples. Matterport3D provides RGBD scans and navigable meshes of real-world indoor scenes with object and room type annotations. It is used extensively for robotic tasks like navigation and 3D scene understanding. It offers very detailed scenes with cluttered rooms. The major challenges associated with it are the smaller number of scenes (90), class imbalance in room types, and noisy meshes. Although Matterport3D is a popular dataset for 3D indoor scene understanding, no other work discusses topological segmentation of scenes for this dataset. Since Matterport3D does not provide annotations for transition regions, we manually annotated all doorways in the dataset. We share this extension to the dataset with the community.
@@
Figure 3: Qualitative comparisons on room segmentation task. Our proposed method is able to handle adverse cases such as regions of low density (introduced by the scanning strategy) in the Stuctured3D dataset, as well as curved walls and clutter as seen in the Matterport3D dataset. Colors are only representative of the instance separation.
Baselines and Results. We primarily compare our model against the state-of-the-art RoomFormer [13] and HEAT [4]. Both are transformer-based methods that fit polygons and output an ordered set of vertices. [4] first detects room corners, and then predicts edges, but no doorways. Results are summarized in Table Table I, where we outperform the average precision of the state-of-the-art RoomFormer by 1.9% on Structured3D and by 20% on Matterport3D. In addition to rooms, RoomFormer also detects transition regions i.e. doorways. Our method registers a 10% higher average precision than the SoTA RoomFormer on doorway detection. To demonstrate the effectiveness of our multi-channel representation (top and bottom occupancy maps), we construct a baseline that learns only on a single-channel density map. As in Table Table I, our ablations exhibit an improvement in both doorway detection and room segmentation tasks.